TinaaaaaA:

IF 10 Questions Equal Five points How many Points is 267??

Florisalreadytaken:

ok, so based on that info, we can create this proportion: \[ \frac{10}{5} \times \frac{267}{x} \] just cross multiply, do the math, and you will have an answer. So you know how to do that?